OpenStudy (anonymous):

An open box is made from a square piece of sheet metal 19 inches on a side by cutting identical squares from the corners and turning up the sides. Express the volume of the box, V, as a function of the length of the side of the square cut from each corner, x.

OpenStudy (anonymous):

So you are basically looking for the volume?

OpenStudy (anonymous):

so is it like this V(x) = x(19-2x)^2

OpenStudy (anonymous):

Well the volume would be 15*15*15 = 3375 I suppose you already got that far?

OpenStudy (anonymous):

oh wait 19 in

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

It would be 19*19*19 = 6859

OpenStudy (anonymous):

my choices are just functions not answers A. V(x) = 361x B. V(x) = (19-2x)^2 C. V(x) = x(19-2x) D. V(x) = x(19-2x)^2

OpenStudy (anonymous):

V = (19-2x)^2*x

OpenStudy (anonymous):

that is what I got but it don't match my choices would it be D

OpenStudy (anonymous):

Yes :D

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

:P